OpenStudy (rainbow_dashie):

The juice dispenser in the cafeteria holds 29 2/3 pints of juice. The glasses in the cafeteria only hold 1/3 of a pint of juice. How many full glasses can be served at breakfast?

OpenStudy (igreen):

Just divide.. 29 2/3 / 1/3 = ? @Rainbow_dashie

OpenStudy (rainbow_dashie):

_ 9.8 is the decimal form

OpenStudy (igreen):

No..check again.

OpenStudy (rainbow_dashie):

30

OpenStudy (igreen):

First we can change \(29 \dfrac{2}{3}\) into an improper fraction. Multiply the denominator by the whole number, then add it to the numerator: \(3 \cdot 29 + 2 = \dfrac{89}{3}\) Now can you divide? \(\dfrac{89}{3} \div \dfrac{1}{3} = \dfrac{89}{3} \cdot 3\) Can you multiply 89/3 * 3? @Rainbow_dashie

OpenStudy (igreen):

30 is wrong too.

OpenStudy (rainbow_dashie):

89

OpenStudy (igreen):

Yep! So that's your answer.

OpenStudy (rainbow_dashie):

thank you!

OpenStudy (rainbow_dashie):

The price of a gallon of gasoline is $4.05. The price when Ryan's mother started driving was 1/9 of the current price. What was the price of gasoline when Ryan's mother started driving?